4) An investment offers to pay $ 100 a year forever starting at the end of year 6. If the interest rate is 8%, then what is the investment's value today? (Take note: first payment happens at the end of 6th year, not the end of 1st year.)

Explain your work
Answer #1

Correct Answer is $850.73

Explanation:

As the investment offers $100 from the end of 6th Year till perpetuity,

The present value at the end of 5th Year = Coupon Amount / Rate = $100 / 0.08 = $1,250

Hence, the present value of $ 1,250 today will be = P / (1+r)5 = $1,250 / (1+0.08)5 = $1,250 / 1.469328 = $850.73